# Longitudinal Abuse and Sentiment Analysis of Hollywood Movie Dialogues using LLMs

### v2

@@ 6 Conclusion
We employed LLMs for a longitudinal study of abuse and sentiment analysis of subtitle dialogues from prominent Hollywood movies spanning seven decades (1950-2024), including Oscar-nominated and top ten blockbuster movies. We refined LLMs using specific sentiment analysis and abusive datasets. to discern significant trends and shifts in the emotional and abusive content within movie dialogues.
@@
Our results reveal a complex landscape of emotional expression in movies, with notable fluctuations in the prevalence of various sentiments and abusive content. Our findings indicate that while positive emotions such as humour and optimism have consistently been integral to movie dialogues, the representation of negative emotions like anger, sadness, and anxiety has varied significantly, often reflecting broader social and cultural changes. Our categorisation of films into four broad genres—action, comedy, drama, and thriller—provided further granularity to our analysis. Each genre exhibited distinct patterns of emotional and abusive content, with overlaps, since the majority of movies can be classified into two categories.
@@
One of the key insights is the gradual increase of abusive content in movie dialogues over time, especially in the last two decades, where Oscar-nominated movies overtook the top ten blockbusters. This trend suggests a gradual shift in the industry’s approach to depicting violence and abuse, potentially influenced by evolving societal norms and regulatory frameworks.
@@ 8 Acknowledgement
We thank UNSW School of Mathematics and Statistics, DATA-9020 Group-H members: Jiaying Zhu,Yuxin Hou, Ziyang Feng, Zimeng Luo for their contributions earlier in the project.

### v3

@@ 6 Conclusions
We employed language models for a longitudinal study of abuse detection and sentiment analysis of subtitle dialogues from prominent Hollywood movies spanning seven decades (1950-2024), including Oscar-nominated and top ten blockbuster movies. We refined language models using specific sentiment analysis and abusive datasets to discern significant trends and shifts in the emotional and abusive content within movie dialogues.
@@
Our results reveal a complex landscape of emotional expression in movies, with notable fluctuations in the prevalence of various sentiments and abusive content. Our findings indicate that while positive emotions such as humour and optimism have consistently been integral to movie dialogues. The representation of negative emotions such as anger, sadness, and anxiety has varied significantly, often reflecting broader social and cultural changes. Our categorisation of films into four broad genres—action, comedy, drama, and thriller—provided further granularity to our analysis. Each genre exhibited distinct patterns of emotional and abusive content, with overlaps, since the majority of movies can be classified into two categories.
@@
One of the key insights is the gradual increase of abusive content in movie dialogues over time, especially in the last two decades, where Oscar-nominated movies overtook the top ten blockbusters. This trend suggests a gradual shift in the industry’s approach to depicting violence and abuse, potentially influenced by evolving societal norms and regulatory frameworks. Our study relies solely on textual subtitles and does not incorporate audio or visual modalities, which are crucial for a more holistic interpretation of sentiment and violence in films. Future work can explore multimodal approaches that integrate speech prosody, facial expressions, and visual context, enabling a more comprehensive analysis of cinematic content.
@@ 8 Acknowledgement
We thank UNSW School of Mathematics and Statistics, DATA-9020 Group-H members: Jiaying Zhu, Yuxin Hou, Ziyang Feng, Zimeng Luo, for their contributions earlier in the project.